OpenStudy (anonymous):

I don't see the reasoning in going straight from this limit to the answer is one.. someone help. (((1+n^(1/7)+n^(1/6))/(2-n+n^(4/3))) * (n^(7/6)) / 1

OpenStudy (anonymous):

\[((1+ \sqrt[7]{n}+\sqrt[6]{n})\div(2-n+ \sqrt[3]{n ^{4}}))(\sqrt[6]{n ^{7}})\] Written out a little nicer

OpenStudy (anonymous):

Do u need to calculate limit?

OpenStudy (anonymous):

Well I'm doing a limit comparison test, and they went straight from this mess to the limit equals one.. I just need to know this limit doesn't equal 0 really..

OpenStudy (anonymous):

And it's as it goes to infinity

OpenStudy (anonymous):

check it out

OpenStudy (anonymous):

I already wolframmed it.. I clicked show steps and there were none, lol. but i'll check it out

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

yeah.. still no steps

OpenStudy (anonymous):

Lim comparison test, they took the function they were evaluating, they took another function that they know how it behaves, function a/function b take lim. Limit is 1is helpful info: lim>0 says function a and function b either both converge or both diverge.

OpenStudy (anonymous):

Yeah, I need help understanding why exactly that limit it one. It looks like there is too much going on to start L'hopitaling it.. And the top and both both have different powers when multiplied through by n^(7/6)

OpenStudy (anonymous):

For better readability: \[\frac{1 + n^{1/7} + n^{1/6}}{2-n +n^{4/3}} * \frac{n^{7/5}}{1}\]

OpenStudy (anonymous):

Is that right?

OpenStudy (anonymous):

Simplify, take lim of n^(highest exponent)/n^(highest exponent)

OpenStudy (anonymous):

yes except it's n^(7/6) / 1 at the end

OpenStudy (anonymous):

Ok, so distribute that n through we get \[\frac{n^{7/6} + n^{1/7 + 5/7} + n^{1/6 + 5/7}}{2-n + n^{4/3}}\]

OpenStudy (anonymous):

Err that's wrong.

OpenStudy (anonymous):

Got my fractions upside-down \[\frac{n^{7/6} + n^{1/7 + 7/6} + n^{1/6 + 7/6}}{2-n+n^{4/3}}\]

OpenStudy (anonymous):

lol.. thats just perfect isn't it

OpenStudy (anonymous):

oops...

OpenStudy (anonymous):

I was trying to multiply the exponents.. I see why it's one now

OpenStudy (anonymous):

So on top we have the largest power of n being \( n^{8/6}\)

OpenStudy (anonymous):

And on bottom, same thing.

OpenStudy (anonymous):

So yeah, 1.

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

When you multiply powers of the same base you add their exponents.

OpenStudy (anonymous):

yeah....